9) A 2 kg object is thrown vertically upward with an initial velocity of 10 m/s. (a) What is its velocity at a point halfway to the maximum height it will attain? (a) What is the total time required for it to return to its initial position? (c) What is its acceleration at the maximum height? (d) What is its average speed during the complete trip? (neglect air resistance
Answer #1

a) u = 10 m/s

max height= 10^2 / ( 2 x 9.8) = 5.102 m

time to reach max height = 10/9.8= 1.02 seconds

half way to max height = 5.102/ 2= 2.551 m

v^2 = 10^2 - 2 (9.8) ( 2.551)

v ^2 = 50

v = 7.07 m/s apprx

b) total time of flight = 2 ( 1.02 seconds)= 2.204 seconds

c) g = -9.8 m/s^2

d) average speed = total distance/ total time=2( 5.012) / 2( 1.02) = 4.914 m/s apprx